Astrid1:

ite ive tried enough 😂 help me here i mightve 4gotten sum stuff

Astrid1:

Sollums:

3

Sollums:

its the 3rd answer

Sollums:

@astrid1 wrote:
did i get it right?

Sollums:

yall typing paragraphs?

Astrid1:

So what I did was \[\frac{ y_2-y_2 }{ x_2-x_1 }= \frac{ 5-4 }{ -2-1 }=- \frac{ 1 }{ 3 }\] Not sure if thats right 😂

Astrid1:

@sollums wrote:
its the 3rd answer
we both got the same answer :\ so im guessing im right?

Sollums:

i use desmos my teacher says we can use it on ohio state test so

Sollums:

@astrid1 wrote:
@sollums wrote:
its the 3rd answer
we both got the same answer :\ so im guessing im right?
....

Astrid1:

Oh, ion use desmos.

Astrid1:

My math teacher likes it but to me its too confusing.

Extrinix:

So the equation is \(\dfrac{y_2-y_1}{x_2-x_1}\). And for this, you plug in the coordinates for your variables. \(\dfrac{4-5}{1-(-2)}\) which gives you \(-\dfrac{1}{3}\) So yes, you're correct on that.

Astrid1:

ey B)

Sollums:

see

Extrinix:

You flip flopped \(x_2\) and \(x_1\) though.

Astrid1:

@sollums wrote:
see
that makes no sense

Astrid1:

@extrinix wrote:
You flip flopped \(x_2\) and \(x_1\) though.
wut did I?

Astrid1:

O lmao I was writing off my notes LMAO I did. I wrote thme backwards.

Extrinix:

@astrid1 wrote:
So what I did was \[\frac{ y_2-y_2 }{ x_2-x_1 }= \frac{ 5-4 }{ -2-1 }=- \frac{ 1 }{ 3 }\] Not sure if thats right 😂
You said \(-2-1\) instead of \(1-(-2)\) Although both would be correct

Extrinix:

Yeah exactly

Astrid1:

yee ty <3

Extrinix:

np <3

Sollums:

@astrid1 wrote:
yee ty <3
yw

Extrinix:

You didn't do anything. You gave a direct answer, which was against the site rules, so sthu.

Astrid1:

Chile bro 😂

Astrid1:

I alr reported him